### v5

# Brain states analysis of EEG predicts multiple sclerosis and mirrors disease duration and burden

Abstract
Background: Any treatment of multiple sclerosis should preserve mental function, considering how cognitive deterioration interferes with quality of life. However, mental assessment is still realized with neuro-psychological tests without monitoring cognition on neurobiological grounds whereas the ongoing neural activity is readily observable and readable.
Objective: The proposed method deciphers electrical brain states which as multi-dimensional cognetoms quantitatively discriminate normal from pathological patterns in an EEG.
Method: Baseline recordings from a prior EEG study of 93 subjects, 37 with MS, were analyzed. Spectral bands served to compute cognetoms and categorize subsequent feature combination sets.
Result: A significant correlation arose between brain states predictors, clinical data and disease duration. Using cognetoms and spectral bands, a cross-sectional comparison separated patients from controls with a precision of 82% while using bands alone arrived at 64%.
Conclusion: Brain states analysis successfully distinguishes controls from patients with MS. The congruity with disease duration is a neurobiological indicator for disease accumulation over time. Our results imply that data-driven comparisons of EEG data may complement customary diagnostic methods in neurology and psychiatry. However, thinking ahead for quantitative monitoring of disease time course and treatment efficacy, we hope to have established the analytic principles applicable to longitudinal clinical studies.
Keywords multiple sclerosis, EEG, brain states analysis, machine learning, monitoring, disease burden
1 Introduction
Mental illness is an enormous burden for societies worldwide (Smith, 2011; Vigo et al., 2016). Above all it is the decay of cognitive functioning, the essence of human reasoning, that turns a mental illness into a devastating and tangible experience for patient and surrounding alike. Nonetheless diagnosis and staging of diseases still rely on customary clinical, laboratory and imaging tests by which cognitive integrity is not assessed on biological grounds. Electrical scalp recordings exist for a century and were used decades ago as diagnostic evidence in multiple sclerosis although the analytic methods lacked the quantitative aspect for relating brain measures to disease time course or treatment outcome. We believe that a modern data-driven functional brain mapping analysis approach using non-invasive EEG data will meaningfully enhance the prognostic power for clinical trials and facilitate the routine assessment of drug efficacy and safety through objective measures for a changing cognito-electrical substrate.
The central nervous system (CNS) is an electrical organ and invokes with each mental act a plethora of neural activities instantly reflected in the electrical brain space and recordable on the skull surface with electrode arrays (Buzsáki, 2005). Electrical states of brain functioning, coined cognetoms, change at any given moment (Koenig et al., 1999; Samdin et al., 2017) and are traced with physiological measures while anatomical brain lesions evolve slower over days or weeks and are monitored with MRI exams (Eshaghi et al., 2021; Tommasin et al., 2021; Lazzarotto et al., 2024). The crucial difference is in the immediacy and temporal accuracy of electrical measures and the dispersed soft demarcation of electrical phenomena on the skull surface that mirror underlying distributed brain activity. These diffuse and perpetually evolving cloud-like arrangements are due to oscillating neural networks and reflect equally housekeeping and cognitive neural processes (Buzsáki, 2005).
MS is a chronic, relapsing-remitting, inflammatory disease with multiple foci of demyelination disseminated over space and time throughout the CNS. It can lead to impaired cognitive reserves and capacity (Maggi et al., 2025) disability (Cobo-Calvo et al., 2025) and early death. Northern countries show a high frequency for MS with the highest prevalence in Canada of 0.3% (Amankwah et al., 2017).
Recent advances in pharmaceutical research for aggressive disease-modifying treatments lead to a surge of novel drugs for MS over the last two decades. Currently, there are 18 disease-modifying therapies approved by Health Canada (relapsing remitting form) while heaps of new compounds are being tested in clinical trials for clinical and MRI outcomes. The desired effect of drug therapy on cognitive preservation (Falet et al., 2022) however is still conventionally assessed with questionnaires and neuro-psychological tests although a recent upswing of publications using functional brain measures in MS is observable (Kisler et al., 2020; Paolicelli et al., 2021; Santinelli et al., 2021; Mohseni and Moghaddasi, 2022; Sorrentino et al., 2022).
We adopted the statistical principles of our former brain states analysis method (Janoos et al., 2011; Mórocz et al., 2012; Janoos et al., 2013) and tested the new method on data taken from a prior EEG study about multiple sclerosis in healthy subjects and patients (Jouzizadeh et al., 2021). Our motivation to analyze such cross-sectional cohort data roots in the current lack of access to serial EEG measures in MS patients. While our methods are primarily developed for the purpose of comparing serial EEG measures longitudinally within the same individual we felt confident that the data modeling phase could also generate robust enough brain state maps from single-time-point data to meaningfully support a cross-sectional comparison of brain states features between two subject cohorts. We show here the tangential application of a brain states EEG analysis method that successfully distinguishes healthy data from MS data.
Figure 1: (A) Result of dimension reduction with un-supervised version of UMAP manifold learning algorithm. The sequence of brain states for each individual data is embedded to a two-dimensional space. Each point corresponds to one individual, red color with MS patients, blue with control subjects. (B) Result of dimension reduction with supervised version of UMAP manifold learning algorithm. The sequence of brain states for each individual data is embedded to a two-dimensional space. Each point corresponds to one individual. We split the data into train (60%) and test (40%) subsets of the data. Green corresponds with the train subset of control subjects and blue with the test subset, then orange with train subset of MS patients and red with the test subset. In both plots, there is no meaning to axis orientation and values, only the relative position of points in two-dimensional space is relevant as well as the density distribution through clustering.
2 Materials & Methods
2.1 Subjects and EEG recording
The EEG data of 93 subjects were taken from an earlier project where participant criteria and recording sessions have been described in detail in the original publication (Jouzizadeh et al., 2021). Participants signed consent forms for all procedures and data analyses in accordance with the Declaration of Helsinki. The anonymized data of 37 subjects diagnosed with MS (24 female, 13 male, mean age 29.5 years, age range 20-53 years) and 56 healthy volunteers (29 female, 27 male, mean age 20.7 years, age range 18-31 years) were taken into consideration for this data analysis project. The mean of the Expanded Disability Status Scale (EDSS) scores for MS patients at EEG acquisition time was 1.96 (EDSS score range 0-7). The diagnosis of MS was based on the revised McDonald criteria for relapsing-remitting MS (RRMS) as used at Boualicina Hospital, Sari, Iran. The mean disease duration after confirmed diagnosis was at EEG recording time 5.1 years (range 0-18 years). Resting-state EEG was recorded for 5 minutes for each condition, eyes-open (EO) and eyes-closed (EC), using a Brainmaster Discovery24 data acquisition system and 19 scalp electrode positions at 256 Hz sampling rate.
2.2 Data preparation
Raw binary sensor data were extracted as time stream vectors from original EEG data files sans further metadata or personal information. Electrode labels and anatomical positional information were discarded and played no further role during data analysis. Attention was paid to maximally preserve the fine-grained intrinsic physiological information present in the data (Meisler et al., 2019) for the brain states generation phase further downstream in the processing pipeline. For these reasons EEG data were taken as is without applying common data filtering, cleansing and artifact removal techniques for noise reduction, faulty electrodes, eye blink, cardiac or muscle activity, and without introducing pre-processing related artifacts (McDermott et al., 2021). This avenue is possible because the data processing methods during brain state detection and later analysis phases intrinsically suppress noise components in the data throughout the entire brain states processing pipeline (Janoos et al., 2011; Mórocz et al., 2012) and render the approach fundamentally apt and robust in dealing with clinical-grade real-life data (Janoos et al., 2013). This ‘automatic noise robustness’ lies in the fact that brain states detected throughout the data sets bear inter-individual characteristics and are hence not prone to include meaningless individual noise properties not universally present across the entire data set.
2.3 Brain states detection (BSD)
We define brain states, or cognetoms, as snapshots of repeatable patterns of highly correlated signals. Brain states are understood as moments of distributed brain activity of a given time point (Koenig et al., 1999; Janoos et al., 2011; Mórocz et al., 2012). In a first move a limited set of eight brain states are computed. These sets serve as lowest-level elements, or level-0 cognetoms, quasi as alphabet letters or building blocks to use in future aggregation cascade models of ever increasing cognetom complexity and abstraction level. EEG recordings were segmented into data epochs where each individual epoch, a brain state, consists of 500 EEG frames (scan frequency at 256Hz, epoch length 1.953sec), and are described by five EEG frequency band amplitudes (delta: 0-4Hz, theta: 4-8Hz, alpha: 8-14Hz, beta : 14-32Hz, gamma: 32-45Hz) that are extracted from raw EEG data.
In a second move the best match distribution of brain states is detected for each time point along the EEG data time axis with weights shown for each brain state. In other words, our approach translates the multi-channel raw EEG signal of each subject into a sequence of brain states as qualitative feature with the associated weights as numerical quantitative feature.
3 Results
3.1 Brain states analysis (BSA)
In order to capture differences between the EEG data of healthy persons and patients with MS two approaches were applied : manifold learning and catboost classifier. For this, 80 distinct feature combinations were built and are described as combinations of 8 brain state weights, two experimental recording conditions (eyes-open and eyes-closed) and five EEG frequency band amplitudes as mentioned.
The first approach, manifold learning, reduces the dimensionality of the individual data (McInnes et al., 2018). The previously prepared sequence of brain states and weights are used to embed the individual data into a two-dimensional space. Each point in that space corresponds with one person. First we applied the non-supervised version of the UMAP algorithm for which the results reveal a distinct trend separating control from MS cases (upper figure 1). Next we applied the supervised version by splitting data into train (60%) and test (40%) subsets of data. The trained model produced the two-dimensional embedding with high separability between the two subject groups (lower figure 1).
Our second approach used the catboost classifier method (Dorogush et al., 2018) which first splits the data into train (60%) and test (40%) subsets of data, then trains the model, and last evaluate the quality of the model on the test data subset. The quality of the model is described by precision and F-score values.
In sum, our very first implementation of the brain states approach for EEG analysis shows a successful performance for detecting MS in common clinical low-resolution EEG data sets at a high precision of 82%.
separation based on precision F-score
frequency bands alone 0.64 0.61
frequency bands and brain states 0.82 0.78
Table 1: Performance of MS detection with catboost classifier. The inclusion of brain states features augments significantly the precision for distinguishing EEG data sets in patients with MS vs healthy subjects as compared to taking into account the frequency band features alone.
3.2 Benefit of the brain states approach
We demonstrate a direct benefit for including brain states features in the analysis as compared to taking into account only the frequency band features typically used for EEG analysis (table 1). A benchmark for the two EEG analysis approaches show an impressive performance boost of 18% for distinguishing EEG data sets of patients with MS versus healthy subjects in favor of the combined method. The precision for detecting MS using frequency bands alone reaches 64% while adding information about brain states arrives at 82%.
Figure 2: Specific feature analysis for MS detection. The relative importance of feature combination (brain states vs frequency bands vs experimental conditions) is shown. (A) The upper figure part depicts ordered feature combinations (x-axis) vs feature importance (y-axis). (B) The lower part shows the comparison between the control vs the patient group regarding feature importance. The blue line stands for the control group, the orange line for MS. Only the 40 top most important feature combinations are shown.
3.3 Specific feature combinations for MS detection
The relative importance of 80 feature combinations (8 brain states x 5 frequency bands x 2 experimental conditions) that separate the two subject groups into healthy vs MS is outlined in figure 2. We observe that a set of certain feature combinations that have a high relative importance, include both eyes-open and eyes-closed conditions and the delta frequency band appear to drive the separation between healthy and MS subjects.
3.4 Disease predictor compared to clinical data
The brain states analysis technique separates with high probability healthy controls from patients with MS (figures 1 and 3). In order to explain this rather facile achievement, considering the data originate from single one-time clinical EEG measurements, we examined additional objective neurological measures such as EDSS values or plain disease duration which we hypothesized will correlate with our EEG analysis. These last two clinical measures arguably reflect disease related accumulation of neurological deficits.
Both plots in figure 3 imply a significant correlation of these two clinical measures with the brain states predictor space, which, we surmise, separates the two subject groups, the more the disease burden in the CNS increases. In fact, all patients to the far left in the ‘control’ brain states predictor space (marked green with letters from a to g) feature a short disease history of less than two years (mean 0.8 years) and cited motor and sensory deficits which may originate from lesions in the spinal cord rather than the upper brain parts. In contrast, the patients to the right (marked red) with longer disease histories (mean 6.0 years), self-reported also vision, balance and brain stem disturbances beside only motor and sensory issues in the limbs. This suggests more complex neurological deficit patterns that are likely related to an increase of lesion load in the cerebrum itself, the de-facto generator for the EEG signal.
Figure 3: Distinction of control subjects from patients with MS through brain states prediction. The disease predictor value corresponds to the rotated coordinates as seen in Fig. 1B. A rotation was performed using the principal components of the point distribution where the predictor value is considered as the primary principal component. (A) The upper graph illustrates this separability plotted against EDSS measures that were obtained in MS. Patients are shown as red or green dots while controls feature blue dots. Seven patients are found as outliers on the left side of the disease predictor axis and are marked with green characters (a-g) for further discussion under Results. A linear regression is shown as gray line. (B) The lower graph plots separability against disease duration in years. Two patients (d,f) lack data on disease duration and hence are not included here. Linear regression is shown as gray trend line.
4 Discussion
We successfully transformed a brain states hypothesis (Koenig et al., 1999; Janoos et al., 2011) into an analytic method for processing EEG data using a quantitative embedding and modeling approach for brain states and applied the new technique to separate the two subject groups of an earlier EEG study about MS (Jouzizadeh et al., 2021). We observed several beneficial aspects when brain states analysis is added in a context as described here.
4.1 Implications
First we demonstrate with an initial implementation of our novel algorithm, a two-phase processing pipeline with brain states detection and brain states analysis, that already regular clinical EEG data clearly carry the necessary information to discriminate two cohorts as shown here. This is to our knowledge a first successful demonstration of such a hypothesis-free space-agnostic data-driven full-brain raw-data EEG analysis approach for MS. Second we state that the inclusion of brain states in the analysis is sensible and significantly enhances the statistical robustness of the approach. Third we believe this brain states approach is easily generalizable to other disease conditions and importantly also for the intra-individual monitoring of serially acquired longitudinal EEG data.
The current method readily distinguished healthy from MS cases using clinical EEG data in spite of obstacles faced with : i) early stage of the software implementation; ii) swap in the original purpose of the approach from monitoring serial data to comparing subjects; iii) low information density in common clinical EEG recordings. We explain the analysis outcome with flow derangement between electrical mental processes in MS secondary to altered physical brain properties (Kisler et al., 2020; Paolicelli et al., 2021). This was newly confirmed in a recent study about disturbances of conductance velocity through white matter lesions (Sorrentino et al., 2022). We currently offer no method to quantify processing speed in brain states but we know already that certain feature combinations of high relevance carry slow delta EEG components associated with both eye-open and eyes-closed conditions that appear to drive the separation of controls from patients (see figure 2).
We speculate that even just small conductance alterations in MS will perceptibly skew the ’electric symphony’ among the many brain activities, ’handshakes’, synchronicity and interplay between resonating loops (Buzsáki, 2005), all the while the patient still compensates without visible clinical deficit. However, our concept about internal clockworks being disturbed between brain activities will remain hypothetical without having access to high-quality EEG data (Formica et al., 2025).
4.2 Significance
That this method succeeds on clinical-grade EEG data rests on two facts, in the processing of complete brain states and the characterizing of entire EEG recordings in form of comprehensive brain states sequences without cleansing noisy epochs. These two points highlight potential advantages of the method in that older less-advanced EEG databases could be sensibly re-analyzed with modern brain states analysis techniques, while collecting new EEG data in MS or other conditions could for budgetary reasons happen on affordable low-resolution EEG recording devices. Having said that, only high sensor density (Formica et al., 2025), high scan frequency and rich cognitive stimulative task conditions will leverage dense enough information in the EEG and hence contribute to deep brain states abstraction models. While clinical EEG data come in their own right only the highest-quality EEG data can deliver the fodder to study the internal syntax inside multi-hierarchical cognetom cascade models.
4.3 Comparison with existing methods
The here used EEG data were analyzed successfully with two methods, an earlier approach (Jouzizadeh et al., 2021) and the currently presented approach, as both separated the data into two groups, control subjects and patients with MS. The fundamentally distinct nature of the two methods, however, makes a comparative explanation for the similar results difficult. The former method relies on differentiating specific features found in space and EEG frequency bands inside heavily cleaned EEG excerpts (Jouzizadeh et al., 2021). The latter method distinguishes features in a space and time independent, non-localizational manner tracing diffusely distributed signal patterns inside brain states that are registered as comprehensive state sequences over the time course of an entire EEG recording. In essence, the here introduced latter brain states method raises the analytic precision when added to the former ‘baseline’ method which solely relies on frequency band features (table 1). It is apt to detect subtle and dispersed changes within the ’clockwork’ of EEG rhythms because these signal modulations are holistically picked up as identifiable and quantifiable biases encoded within brain states sequences secondary to disease related alterations in electrical brain rhythms.
The pace for new publications using EEG to monitor MS with modern methods is picking up. It includes approaches such as describing changes in brain dynamics (Gschwind et al., 2016) in microstates (Schependom et al., 2019). using quantitative EEG with frequency power analysis (Classen et al., 2019) for chronic neuropathic pain (Kisler et al., 2020), obstacle avoidance during gait (Santinelli et al., 2021), after optic neuritis (Pawlitzki et al., 2020), in gray matter lesions (Salim et al., 2021), tracking in vivo functional remyelination (Farley et al., 2019) or cognitive impairment (Kiiski et al., 2018; Paolicelli et al., 2021; Sarrias-Arrabal et al., 2021; Akbarian et al., 2024; Vlieger et al., 2024) or fatigue (Baldini et al., 2024; Pauletti et al., 2025) or motor disorder (Bardel et al., 2024a; Leodori et al., 2024) or auditory processing (Bissonnette et al., 2023), measuring event-related desynchronization latency (Carrere et al., 2025; Martell et al., 2025), advanced analysis methods (Mohseni and Moghaddasi, 2022; Şaşmaz Karacan and Saraoğlu, 2024; Aslan and Akşahin, 2025) with MEG data (Costers et al., 2021), resting state EEG in Alzheimer’s disease (Babiloni et al., 2021), in cognitive workload studies of neurodegenerative diseases (Figueroa-Vargas et al., 2020; Cano et al., 2024) and finally literature reviews about the use neurophysiological methods in MS (Hernandez et al., 2023; Bardel et al., 2024b; Cruciani et al., 2024; Ferreira et al., 2024; Formica et al., 2025).
The central distinction, however, of the here introduced brain states method versus those referenced above is that its modeling approach, though at an early state of development and implementation, suits at once a full range of features extracted during the data preparation phase, and operates in that sense by default much more comprehensively than techniques that focus on one specific signal feature.
4.4 Limitations
First the low spatial and temporal resolution of the analyzed EEG data, albeit acquired at clinical grade, is an important limiting factor whereas instead an increase in those acquisition parameters is indeed requisite for arriving at the maximal potential of the presented analysis technique. Second the two non-engaging experimental conditions, eyes-closed and eyes-open, further limit what instead a more engaging cognitive session could offer through exposing participants to standard audio-visual stimuli or to disease-specific experimental tasks which we expect will greatly improve sensitivity of the approach to subject- or disease-specific patterns in the EEG signal. Third the brain states epoch window kept for data-technical reasons at two seconds width proved to work surprisingly well but does no justice to the speed ephemeral electrical brain activities operate at. In other words, it is the acquisition of high-quality EEG data that will crucially counter such limitations and properly account for the short-lived transient nature of informative high-frequency signal elements.
4.5 Future objectives
We see the purpose of this publication in first line in introducing a novel hypothesis-free data-driven analysis method tested on EEG jointly with clinical data of patients with MS. However another raison-d’être is our desire to raise interest in the community towards functional brain data in longitudinal clinical studies in MS in the form of serial EEG measures (Classen et al., 2019; Farley et al., 2019; Schependom et al., 2019; Figueroa-Vargas et al., 2020; Jouzizadeh et al., 2021). Moreover we expect for the future an MS-specific EEG repository to foster novel quantitative neurocognitive methods on disease progression (Tur et al., 2023) and drug efficacy and to parallel in how the International Progressive MS Alliance (IPMSA) (Thompson et al., 2022) catalyzes research between academia and pharmaceutical industry for MRI data in MS (Eshaghi et al., 2021; Falet et al., 2022).
5 Conclusion
We introduced a novel hypothesis-free quantitative full-brain analysis approach and successfully separated EEG data paired with clinical evidence of patients suffering from MS from those of healthy subjects. Thinking forward, we envision quantitative cognitive imaging using serial EEG data to become a future main player for longitudinal clinical studies where functional neurobiological outcome measures are key, be that for monitoring MS or other illnesses with prolonged time courses such as Alzheimer, Parkinson or Huntington disease, schizophrenia, depression, migraine, HIV or systemic chemotherapy. However as the knowledge base with cognitive imaging techniques grows we believe that EEG will, as in the here presented case, also have a say in the diagnosis, subtyping and staging of primary and secondary neurological diseases.
6 Declaration of conflicting interests
The authors declared the following potential conflicts of interest with respect to the research, authorship and publication of this article: István Mórocz and Andrei Mogoutov co-founded Noisis Inc., an early-stage startup.
7 Funding
The authors did not receive any financial support nor any specific grant from funding agencies in the public, commercial, or not-for-profit sectors for the research, authorship, and publication of this article.
8 Data availability statement
Data sharing may be applicable upon request to the data controller R.K.