### v8

# Brain states analysis of EEG predicts multiple sclerosis and mirrors disease duration and burden

Abstract
Background: Any treatment of multiple sclerosis should preserve mental function, considering how cognitive deterioration interferes with quality of life. However, mental assessment is still realized with neuro-psychological tests without monitoring cognition on neurobiological grounds whereas the ongoing neural activity is readily observable and readable.
Objective: The proposed method deciphers electrical brain states which as multi-dimensional cognetoms quantitatively discriminate normal from pathological patterns in an EEG.
Method: Baseline recordings from a prior EEG study of 88 subjects, 36 with MS, were analyzed. Spectral bands served to compute cognetoms and categorize subsequent feature combination sets.
Result: The brain states predictor correlates with disease burden and duration. Using cognetoms and spectral bands, a cross-sectional comparison separated patients from controls with a precision of 85% while using bands alone arrived at 79%.
Conclusion: We demonstrate the efficiency of the quantitative data-driven method based on brain states analysis by contrasting EEG data of patients with MS and healthy subjects. The congruity with disease severity and duration is a neurophysiological indicator for disease accumulation over time. We discuss potential applications of the approach for the monitoring of disease time course and treatment efficacy in longitudinal clinical studies in psychiatry and neurology.
Keywords multiple sclerosis, EEG, brain states analysis, machine learning, monitoring, disease burden
1 Introduction
Mental illness is an enormous burden for societies worldwide (Smith, 2011; Vigo et al., 2016). Above all it is the decay of cognitive functioning, the essence of human reasoning, that turns a mental illness into a devastating and tangible experience for patient and surrounding alike. Nonetheless diagnosis and staging of diseases still rely on customary clinical, laboratory and imaging tests by which cognitive integrity is not assessed on biological grounds. Electrical scalp recordings have existed for a century and have been used decades ago as diagnostic evidence in multiple sclerosis, although the analytic methods lacked the quantitative aspect needed to relate brain measures to disease time course or treatment outcome. We believe that a modern data-driven functional brain mapping analysis approach using non-invasive EEG data will meaningfully enhance the prognostic power for clinical trials and facilitate the routine assessment of drug efficacy and safety through objective measures for a changing cognito-electrical substrate.
The central nervous system (CNS) is an electrical organ and invokes with each mental act a plethora of neural activities instantly reflected in the electrical brain space and recordable on the skull surface with electrode arrays (Buzsáki, 2005). Electrical states of brain functioning, coined cognetoms, change at any given moment (Koenig et al., 1999; Samdin et al., 2017) and are traced with physiological measures while anatomical brain lesions evolve slower over days or weeks and are monitored with MRI exams (Eshaghi et al., 2021; Tommasin et al., 2021; Lazzarotto et al., 2024; Goebl et al., 2025). The crucial difference is in the immediacy and temporal accuracy of electrical measures and the dispersed soft demarcation of electrical phenomena on the skull surface that mirror underlying distributed brain activity. These diffuse and perpetually evolving cloud-like arrangements are due to oscillating neural networks and reflect equally housekeeping and cognitive neural processes (Buzsáki, 2005).
MS is a chronic, often relapsing‑remitting, inflammatory disease with multiple foci of demyelination disseminated over space and time throughout the CNS. It can lead to impaired cognitive reserves and capacity (Maggi et al., 2025; Pauletti et al., 2025) disability (Cobo-Calvo et al., 2025) and early death. Northern countries show a high frequency for MS, with the highest prevalence in Canada of 0.3% (Amankwah et al., 2017).
Recent advances in pharmaceutical research for aggressive disease-modifying treatments lead to a surge of novel drugs for MS over the last two decades. Currently, there are 18 disease-modifying therapies approved by Health Canada (relapsing remitting form), while heaps of new compounds are being tested in clinical trials for clinical and MRI outcomes. The desired effect of drug therapy on cognitive preservation (Falet et al., 2022), however, is still conventionally assessed with questionnaires and neuro-psychological tests, although a recent upswing of publications using functional brain measures in MS is observable (Kisler et al., 2020; Paolicelli et al., 2021; Santinelli et al., 2021; Sarrias-Arrabal et al., 2021; Mohseni and Moghaddasi, 2022; Sorrentino et al., 2022).
For this work we propose an extension for the concept of brain states analysis (Janoos et al., 2011; Mórocz et al., 2012; Janoos et al., 2013). Instead of the statistical principles based on hidden Markov model (HMM) and functional MRI data we developed a novel method for brain states detection and analysis of EEG signals taken from a prior cross-sectional EEG study about patients with multiple sclerosis and healthy subjects (Jouzizadeh et al., 2021).
Figure 1: (A) Result of dimension reduction with un-supervised version of UMAP manifold learning algorithm. The sequence of brain states for each individual data is embedded to a two-dimensional space. Each point corresponds to one individual, red color with MS patients, blue with control subjects. (B) Result of dimension reduction with supervised version of UMAP manifold learning algorithm. The sequence of brain states for each individual data is embedded to a two-dimensional space. Each point corresponds to one individual. Green corresponds with the train subset of control subjects and blue with the test subset, then yellow with train subset of MS patients and red with the test subset. In both plots, there is no meaning to axis orientation and values, only the relative position of points in two-dimensional space is relevant as well as the density distribution through clustering.
2 Materials & Methods
2.1 Subjects and EEG recording
The EEG data of 88 subjects were taken from an earlier project. Participant criteria and recording sessions have been described in detail in the original publication (Jouzizadeh et al., 2021). Participants signed consent forms for all procedures and data analyses in accordance with the Declaration of Helsinki. The anonymized data of 36 subjects diagnosed with MS (23 female, 13 male, mean age (29.6 years, age range 20-53 years) and 52 healthy volunteers (25 female, 27 male, mean age 20.7 years, age range 18-31 years) were taken into consideration for this data analysis project (see tables in A and B for subject information). The mean of the Expanded Disability Status Scale (EDSS) scores for MS patients at EEG acquisition time was 1.96 (EDSS score range 0-7). The diagnosis of MS was based on the revised McDonald criteria for relapsing-remitting MS (RRMS) as used at Boualicina Hospital, Sari, Iran. The mean disease duration after confirmed diagnosis was at EEG recording time 5.1 years (range 0-18 years). Resting-state EEG was recorded for 5 minutes for each condition, eyes-open (EO) and eyes-closed (EC), using a Brainmaster Discovery24 data acquisition system and 19 scalp electrode positions at 256 Hz sampling rate.
2.2 Data preparation
Raw binary sensor data were extracted as time stream vectors from original EEG data files without further metadata or personal information. Electrode labels and anatomical positional information played no role during data analysis. Attention was paid to maximally preserve the intrinsic physiological information present in the data (Delorme, 2023) for the brain states generation phase further downstream in the processing pipeline. EEG data were taken as is without applying common manual or automated data filtering, cleansing and artifact removal techniques for noise reduction, faulty electrodes, eye blink, cardiac or muscle activity, and hence without introducing data pre-processing related artifacts.
2.3 Brain states detection (BSD)
We were inspired by the similarity of the brain states analysys method we had developed for functional MRI data (Janoos et al., 2011; Mórocz et al., 2012) and the concept of microstates analysis for EEG data introduced by Lehmann et al (Lehmann et al., 1987). Both approaches aim to transform the complex original data into a sequence of discrete, quasi-stable meaningful patterns, also called states. It was shown that the distribution of these states may play a biomarker role in various psychiatric (Koenig et al., 1999; Janoos et al., 2013) and neurological conditions including MS (Gschwind et al., 2016). In the current work, however, a brain state is a typical configuration of spatial activity without actual knowledge of spatial coordinates of electrodes.
We earlier defined brain states, or cognetoms, as snapshots of repeatable patterns of highly correlated signals also understood as moments of distributed brain activity of a given time point (Janoos et al., 2011; Mórocz et al., 2012). In other words, states are defined as patterns of statistical dependencies or spatio-temporal correlations between brain regions measured across EEG electrodes. The states serve as lowest-level elements, or level-0 cognetoms, quasi as alphabet letters or building blocks to use in future aggregation cascade models of ever increasing cognetom complexity and abstraction level.
In a first move a limited set of eight brain states are computed. Raw EEG signals from each channel are transformed into sequences of overlapping spectrograms using the Fast Fourier Transform (FFT) implementation from the cusignal.spectrogram library (Thompson and Nicely, 2021). The recordings are segmented into data epochs where each individual window, a brain state, consists of 500 EEG frames (scan frequency at 256Hz, epoch length 1.953sec), and are described by five EEG frequency band amplitudes (delta: 0-4Hz, theta: 4-8Hz, alpha: 8-14Hz, beta : 14-32Hz, gamma: 32-45Hz). While window size and number of states are varied as technical optimization against detection performance, its length of 2s approximates the scan acquisition times needed for a single multi-slice T2* weighted brain shot typical for functional MRI studies (Mórocz et al., 2012; Janoos et al., 2013). To reduce dimensionality and capture coherent spatial patterns, we grouped EEG channels based on the similarity of their time series. Pairwise distances between channels were computed, followed by hierarchical clustering to identify groups of similar channels. For the feature vector construction, we computed at each time step the mean signal intensity across predefined frequency bands within each cluster of grouped channels. This resulted in a concatenated vector representation summarizing the spectral profile of the EEG activity at each moment. For the identification of brain states, we applied k-means clustering to the resulting vectors to identify a set of prototypical channel-spectral configurations, which we interpret as data-driven representations of transient brain states.
In a second move, for the temporal coding of brain states, or brain states sequence, the best match distribution of brain states is detected for each time point along the EEG data time axis with weights shown for each brain state. Each time step in the original sequence is assigned to its nearest brain state cluster, producing a simplified temporal sequence of discrete brain state codes, each associated with signal intensities across frequency bands. In other words, our approach translates the multi-channel raw EEG signal of each subject into a sequence of brain states as qualitative feature with the associated weights as numerical quantitative feature.
2.4 Brain states analysis (BSA)
In order to capture differences between the EEG data of healthy persons and patients with MS two approaches were applied : manifold learning and catboost classifier. For the subject-level feature extraction, 80 distinct feature combinations were built and are described as combinations of 8 brain state weights, two experimental recording conditions (eyes-open and eyes-closed) and five EEG frequency band amplitudes as mentioned. This compressed representation enabled further statistical analysis. For each subject, we computed distributions over brain states and average signal intensities per state. We used a fixed-length feature vector for each subject, defined by the combination of brain states (8), frequency bands (5), and experimental conditions (2), for the downstream comparative analyses.
The first approach, manifold learning, reduces the dimensionality of the individual data (McInnes et al., 2018). The previously prepared sequence of brain states and weights are used to embed the individual data into a two-dimensional space. Each point in that space corresponds with one person. First we applied the non-supervised version of the UMAP algorithm for which the results reveal a distinct trend separating control from MS cases (upper figure 1). Next we applied the supervised version by splitting the data into train (60%) and test (40%) subsets. The trained model produced the two-dimensional embedding with high separability between the two subject cohorts (lower figure 1).
Our second approach used the catboost classifier model (Dorogush et al., 2018) which applied a 30-fold cross-validation to split the data into train (60%) and test (40%) subsets, then trains the model, and last evaluate the quality of the model on the test data subset. Only the binary labels for MS and non-MS were used for model training. All other clinical or demographic information was taken into account only for the post-analysis procedures. The quality of the model is described by precision, recall, f1-score and support (see table 1).
3 Results
3.1 Benefit of the brain states approach
Our very first implementation of the brain states approach for EEG analysis shows a successful performance for detecting MS in common clinical low-resolution EEG data sets at a high precision of 85%. We demonstrate a direct benefit for including brain states features in the analysis as compared to taking into account only the frequency band features typically used for EEG analysis (table 1). A benchmark for the two EEG analysis approaches show a performance boost of 6% for distinguishing EEG data sets of patients with MS versus healthy subjects in favor of the combined method. The precision for detecting MS using frequency bands alone reaches 79% while adding information about brain states arrives at 85%.
separation based on precision recall f1-score support
frequency bands alone 0.79 0.78 0.78 36
frequency bands & brain states 0.85 0.84 0.84 36
Table 1: Performance of MS detection with catboost classifier. We compare the classifier performance for two scenarios : first the features are defined as frequency bands and experimental conditions such as eyes open and eyes closed. In the second scenario the features are defined as a combination of brain states, frequency bands and experimental conditions. We applied a 30-fold cross-validation to split the data into train (60%) and test (40%) subsets of the data. Taking into account the brain states augments the precision, recall and f1-score for distinguishing EEG data sets in patients with MS vs healthy subjects - as compared to taking into account the frequency band features alone.
3.2 Specific feature combinations for MS detection
The relative importance of 80 feature combinations (8 brain states x 5 frequency bands x 2 experimental conditions) that separate the two cohorts into healthy vs MS is outlined in figure 2. To reduce the risk of overfitting only the first 15 feature combinations of high importance including both eyes-open and eyes-closed conditions and the delta frequency band were selected that appear to drive the separation between healthy and MS subjects. Averaged frequency spectra for both cohorts are shown as graphs in C.
Figure 2: Specific feature analysis for MS detection. The relative importance of feature combination (brain states vs frequency bands vs experimental conditions) is shown. For example the first feature combination is called ‘Delta_5_ec’ which stands for ‘Delta’ frequency band, brain states no. ‘5’, and ‘ec’ for the eyes closed condition. (A) The upper figure part depicts ordered feature combinations (x-axis) vs feature importance (y-axis). (B) The lower part shows the comparison between the control vs the patient cohort regarding feature importance. The blue line stands for the control group, the orange line for MS. Only the 40 top most important feature combinations are shown.
3.3 Disease predictor compared to clinical data
The brain states analysis technique separates with surprising ease and high probability healthy controls from patients with MS (figures 1 and 3). Considering the data to originate from single clinical-grade EEG recordings, we therefore examined additional objective neurological measures such as age, EDSS values or plain disease duration which we hypothesized will correlate with our EEG analysis. Age itself does not indicate an observable effect on MS prediction in both cohorts (top graph A in figure 3). However the next two clinical measures arguably reflect disease related accumulation of neurological deficits. The next three plots in figure 3 imply a correlation of these two clinical measures with the brain states predictor space, which, we surmise, distinguishes the better the two cohorts, the more disease burden over time accumulates in the CNS. Patients with symptoms other than pure motor symptoms show a subtle trend with EDSS computed on 26 such cases (third graph C). Regarding disease duration, however, all patients to the far left in the ‘control’ brain states predictor space (marked green with letters from a to g) feature a short disease duration of less than two years (mean 0.8 years) and cited motor and sensory deficits which may originate from lesions in the spinal cord or lower brain rather than upper brain parts. In contrast, patients to the right marked red, with a mean disease duration of 6.0 years, reported also vision, balance or brain stem disturbances beside motor and sensory issues in the limbs. Still, those patients in red to the right with a brief disease duration of 3 years or less, may in fact, we hypothesize, suffer from a longer, hidden disease duration, albeit clinically compensated, to amass enough supratentorial pathology that finally deranges their EEG patterns. We assume that an increase in lesion load in the cerebrum, the principal generator for the EEG signal, relates just vaguely and non-linearly to the emerging complex clinical neurological deficits, be that in cognition, EDSS, or then in form of clinical disease duration, or ‘true’ disease onset, an unknown entity for that matter.
Figure 3: Distinction of control subjects from MS patients through brain states prediction. Disease predictor values correspond to the rotated coordinates as seen in Fig.1B using principal components of the point distribution with the predictor value considered as primary principal component. (A) The top graph illustrates age distribution as a function of the MS predictor scale. It indicates no observable effect of age on MS prediction. The control outliers on the right side (false positives in blue) do not have higher ages than the controls on the left side (true negatives). Conversely, the MS cases on the left side (false negatives in green) do not have a lower age distribution than the true positives on the right side. (B) The second graph illustrates this separability in the MS cohort plotted against EDSS measures. Patients are shown as red or green dots as in the top graph. The seven outliers on the left are marked with green characters (a-g) for further discussion under Results. (C) The third graph plots separability against EDSS measures only for MS patients having symptoms other then motor symptoms. For these 26 patients a linear regression is shown as gray line. (D) The bottom graph plots separability against disease duration in years. Six patients (including d,f) are excluded for no data on disease duration. Linear regression is shown as gray trend line. The source for the post-analysis correlation calculation is in the values for the brain state predictor space and EDSS or disease duration. The open-source python module linregress from scipy served for calculation (Virtanen et al., 2020).
4 Discussion
We successfully transformed a brain states hypothesis (Janoos et al., 2011) into a new analytic method for processing EEG data using a quantitative embedding and modeling approach for brain states and applied the new technique to separate the two subject cohorts of an earlier EEG study about MS (Jouzizadeh et al., 2021). We observed several beneficial aspects when brain states analysis is added in a context as described here.
4.1 Implications
First we demonstrate with an initial implementation of our novel algorithm, a two-phase processing pipeline with brain states detection and brain states analysis, that already regular clinical EEG data clearly carry the necessary information to discriminate two cohorts as shown here. This is to our knowledge a first successful demonstration of such a hypothesis-free data-driven full-brain raw-data EEG analysis approach for MS.
Second we state that the inclusion of brain states in the analysis is sensible and enhances the statistical robustness of the approach.
Third we believe this brain states approach is easily generalizable to other disease conditions and importantly also for the intra-individual monitoring of serially acquired longitudinal EEG data.
The current method readily distinguished healthy from MS cases using clinical EEG data in spite of obstacles faced with : i) early stage of the software implementation; ii) swap in the original purpose of the approach from monitoring serial data to cross-sectional comparing of subjects; iii) low information density in common clinical EEG recordings. We explain the analysis outcome with flow derangement between electrical mental processes in MS secondary to altered physical brain properties (Kisler et al., 2020; Paolicelli et al., 2021; Salim et al., 2021; Wątorek et al., 2024). This was newly confirmed in recent studies about disturbances of conductance velocity (Sorrentino et al., 2022; Wagner et al., 2025) although we currently offer no method to quantify processing speed in brain states. However, we already know that certain feature combinations of high relevance carry slow EEG components associated with both eye-open and eyes-closed conditions, that appear to prominently drive, but still as part of the entire train of feature combinations, the separation of controls from patients (see figure 2).
We speculate that even just small conductance alterations in MS will perceptibly skew the ’electric symphony’ among the many brain activities, ’handshakes’, synchronicity and interplay between resonating loops (Buzsáki, 2005; Wątorek et al., 2024), all the while the patient still compensates without visible clinical deficit. However, our concept about internal clockworks being disturbed between brain activities will remain hypothetical without having access to high-quality EEG data (Formica et al., 2025; Wątorek et al., 2024).
4.2 Significance
That this method succeeds on clinical-grade EEG data rests on several facts, in the processing of complete brain states, in the intrinsic noise suppression capacity, and in the characterizing of entire EEG recordings in form of comprehensive brain states sequences without cleansing noisy epochs. The need for artifact removal is an ongoing debate in the field while retaining the data in the most original state is a stringent reality (Delorme, 2023). In fact, the removal of artifactual, paradigm-related, non-brain-derived EEG components can detrimentally affect analysis of brain physiological data (McDermott et al., 2021), just as noise filtering may lead to undesired outcomes with poorly characterized noise patterns that are loosely related to behavior (Meisler et al., 2019). Experimental paradigms, however, come necessarily with random patterns of incidental and intentional muscular activities which surface as subject specific chance noise in space and time. The data processing pipeline during the brain state detection and the later analysis phases intrinsically suppress noise components (Janoos et al., 2011; Mórocz et al., 2012) and fundamentally render the approach apt to handle clinical-grade real-life data (Janoos et al., 2013). This ‘design-specific noise robustness’ lies in the fact that brain states are generated for the entire data set and hence hold inter-individual, universally present features but suppress intra-individual periodic or permanent signal properties. This way do patient-specific EEG artifacts due to motor, sensory and other effects not survive in brain states nor lead to subgroups in the analysis nor bias the subsequently computed disease predictor space.
Together these points highlight potential advantages of the method in that older less-advanced EEG databases could be sensibly re-analyzed with modern brain states analysis techniques, while collecting new EEG data in MS or other conditions could for budgetary reasons happen on affordable low-resolution EEG recording devices.
Having said that, only high sensor density (Formica et al., 2025; Wątorek et al., 2024), high scan frequency and rich cognitive stimulative task conditions will leverage dense enough information in the EEG and hence contribute to deep brain states abstraction models. While clinical EEG data come in their own right only the highest-quality EEG data can deliver the fodder to study the internal syntax inside multi-hierarchical cognetom cascade models.
4.3 Limitations
First, the low spatial and temporal resolution of the analyzed EEG data, albeit acquired at clinical grade, is an important limiting factor whereas instead an increase in those acquisition parameters is indeed requisite for arriving at the maximal potential of the presented analysis technique.
Second, the two non-engaging experimental conditions, eyes-closed and eyes-open, further limit what a more engaging cognitive session could offer. We expect that exposing participants to standard audio-visual stimuli or to disease-specific experimental tasks will greatly improve sensitivity of the approach to subject- or disease-specific patterns in the EEG signal.
Third, the difference of mean age between the two cohorts amounts to nine years. We are, however, for the current retrospective analysis not aware of a correlation between age and disease severity (EDSS or symptoms) or disease duration. All subjects still belong to a life stage where we expect task-less experiments to recruit similar brain functions and areas and hence to not age-specifically bias the brain states predictor scale.
Fourth, the brain states epoch window kept for data-technical reasons at two seconds width proved to work surprisingly well but does no justice to the speed at which ephemeral electrical brain activities operate. Hence this window size will be optimized in future iterations of this project with a multi-STFT approach (Ali et al., 2022).
Fifth, the current implementation of the BSD method is not suited for the topographic visualization of the 8 brain states, as it is in the case of the microstates method that follows a clear anatomical preponderance (Lehmann et al., 1987). The spatial information was for code-technical and computational performance optimization reasons lost. Our current method focuses on a channel-spectral approach for analyzing EEG data of patients with MS.
Sixth, we are unaware of the medication given to study participants nor was this information conveyed prior (Jouzizadeh et al., 2021). Certain drug classes are known to modify physiological brain responses where patients with MS may indeed receive periodically benzodiazepines (Van Schependom et al., 2019) for example to control muscle spasms and sleep disturbances, though such symptoms tend to occur at later disease stages. This makes a benzodiazepine induced bias in the analysis outcome of the present study less likely in view of the short mean disease duration of 5.1 years. Similarly does the predominant delta character of the leading brain states feature combinations (see figure 2) point away from the beta frequency presence in benzodiazepine treated patients with MS (Van Schependom et al., 2019).
4.4 Future objectives
First we see the purpose of this publication in the introduction of a novel hypothesis-free data-driven analysis method tested on EEG and clinical data of patients with MS. However another reason is our desire to raise interest in the community towards functional brain data in longitudinal clinical studies in MS in the form of serial EEG measures (Kiiski et al., 2018).
We envision further improvements in the brain states detection (BSD) process through widening the spectrum of methods that analyze the EEG signal. For example the analysis through multifractality and persistence effectively characterizes EEG recordings of patients with MS during eyes open and eyes closed conditions (Wątorek et al., 2024). We consider multifractality and information about long-range temporal correlations, wavelet analysis and microstates, as further procedural candidates that enrich the brain states feature space and enhance model performance for time series analyses in future iterations of this project.
Last, we hope a future MS-specific EEG repository will foster novel quantitative neurocognitive methods on disease progression and drug efficacy (Tur et al., 2023; Montobbio et al., 2025), similar to how the International Progressive MS Alliance (IPMSA) (Thompson et al., 2022) catalyzes research between academia and pharmaceutical industry for MRI data in MS (Eshaghi et al., 2021; Falet et al., 2022).
5 Conclusion
We introduced a novel hypothesis-free quantitative full-brain analysis approach and successfully separated EEG data paired with clinical evidence of patients suffering from MS from those of healthy subjects. Thinking forward, we envision quantitative cognitive imaging using serial EEG data to become a future main player for longitudinal clinical studies where functional neurobiological outcome measures are key, be that for monitoring MS or other neurological and psychiatric illnesses with prolonged time courses. However, as the knowledge base with cognitive imaging techniques grows, we believe that EEG will also play a role in the diagnosis, subtyping and staging of primary and secondary brain diseases.
6 Declaration of conflicting interests
The authors declared the following potential conflicts of interest with respect to the research, authorship and publication of this article: István Mórocz and Andrei Mogoutov co-founded Noisis Inc., an early-stage startup.
7 Funding
The authors did not receive any financial support nor any specific grant from funding agencies in the public, commercial, or not-for-profit sectors for the research, authorship, and publication of this article.
8 Data availability statement
Data sharing may be applicable upon request to the data controller R.K.
Appendix A Table of study participants (control subjects)
index cohort age sex
0 CT 21 m
1 CT 21 m
2 CT 19 m
3 CT 29 f
4 CT 19 f
5 CT 20 f
6 CT 18 f
7 CT 19 m
8 CT 18 m
9 CT 21 m
10 CT 19 m
11 CT 19 f
12 CT 21 f
13 CT 20 f
14 CT 25 f
15 CT 18 m
16 CT 18 m
17 CT 21 m
18 CT 21 m
19 CT 18 m
20 CT 19 m
21 CT 18 f
22 CT 28 m
23 CT 19 f
24 CT 18 f
25 CT 18 f
26 CT 20 f
27 CT 18 f
28 CT 20 f
29 CT 18 f
30 CT 18 m
31 CT 21 f
32 CT 19 f
33 CT 19 f
34 CT 19 m
35 CT 20 m
36 CT 19 m
37 CT 21 m
38 CT 18 m
39 CT 20 m
40 CT 27 m
41 CT 23 m
42 CT 25 f
43 CT 19 f
44 CT 19 m
45 CT 21 m
46 CT 31 m
47 CT 28 f
48 CT 25 f
49 CT 19 m
50 CT 23 f
51 CT 19 f
Table 2: Table is continued with patient data on the next page.
Appendix B Table of study participants (patients with MS)
index cohort age sex disease duration EDSS symptoms code
0 MS 32 f 3 0 bs,vision,sensory
1 MS 28 f 11 7 sensory
2 MS 24 f 0 0 bs,balance,motor b
3 MS 37 f 2 1 sensory
4 MS 40 f 14 0 motor,sensory
5 MS 21 f 2 1 sensory g
6 MS 23 m 1 0 vision
7 MS 31 m 6 0 bs,sensory,motor
8 MS 29 m 5 1.5 bs
9 MS 27 m 4 2.5 motor
10 MS 27 m 0 f
11 MS 32 f 1 0 sensory,motor e
12 MS 26 f 0 7 motor a
13 MS 33 m 0
14 MS 31 f 0
15 MS 26 f 6 2 sensory,motor
16 MS 30 f 0 d
17 MS 20 f 1 4 motor c
18 MS 24 m 0 1 vision
19 MS 21 m 0
20 MS 31 m 0 2 sensory,motor
21 MS 37 m 15 4 vision
22 MS 33 m 5 4.5 vision,motor
23 MS 21 m 1 0 vision,balance
24 MS 53 m 10 6 sensory,motor
25 MS 35 f 8 5.5 sensory,motor
26 MS 40 f 3 2 vision,sensory,motor
27 MS 32 f 5 6 motor
28 MS 28 f 5 0 vision,motor,sensory
29 MS 32 f 3 0 vision,balance
30 MS 30 f 2 1 sensory
31 MS 27 f 13 4.5 vision,sensory
32 MS 26 f 0
33 MS 23 f 1 0 sensory
34 MS 32 f 6 0 sensory,motor
35 MS 25 f 2 3.5 vision,motor
Table 3: List of study participants with multiple sclerosis (MS). The ‘symptoms’ column reports neurological deficits for those patients where such clinical information was available. The related notes in the source data were of sparse and moderate quality. For simplicity, the neurological deficits were binned into 5 categories such as [ balance, bs, motor, sensory, vision ] (with ‘bs’ for brain stem). The ‘code’ column with alphabetic characters (a-g) stands for the seven patients that were found as outliers on the left side of the disease predictor axis marked in green letters (a-g) in figure 3.
Appendix C Mean spectra for EEG frequency bands
Figure 4: Graph showing mean spectra for EEG frequency bands. The average for each of the five common EEG frequency bands ($\delta$, $\theta$, $\alpha$, $\beta$ and $\gamma$) was computed across all subjects, separated for control and MS cohorts, and for the two experimental conditions. The eyes-open (EO) condition shows bars at full opacity, and the eyes-closed (EC) condition at half opacity.
[…]